# Concept-based Adversarial Attack: a Probabilistic Perspective

### v1

@@ Abstract
We propose a concept-based adversarial attack framework that extends beyond single-image perturbations by adopting a probabilistic perspective. Rather than modifying a single image, our method operates on an entire concept — represented by a probabilistic generative model or a set of images — to generate diverse adversarial examples. Preserving the concept is essential, as it ensures that the resulting adversarial images remain identifiable as instances of the original underlying category or identity. By sampling from this concept-based adversarial distribution, we generate images that maintain the original concept but vary in pose, viewpoint, or background, thereby misleading the classifier. Mathematically, this framework remains consistent with traditional adversarial attacks in a principled manner. Our theoretical and empirical results demonstrate that concept-based adversarial attacks yield more diverse adversarial examples and effectively preserve the underlying concept, while achieving higher attack efficiency.
@@ 1 Introduction
Adversarial attacks aim to deceive a classifier while preserving the original meaning of the input object [3, 9, 24, 25]. We refer to the manipulated instance as an adversarial example. Early work by Szegedy et al. [40] introduced adversarial attacks against deep learning models for images.
@@
In the image-based adversarial setting, it is widely accepted that controlling the geometric distance between the adversarial example and its original image is crucial for maintaining the original image’s meaning. Consequently, many adversarial attack algorithms constrain perturbations using norms such as $L_{1}$, $L_{2}$, or $L_{\infty}$. Moreover, given the rapid progress in machine learning research, fair comparisons across different adversarial methods have become essential. Numerous benchmarks and competitions [8, 26] therefore focus on attack success rates under the constraint that geometric distance does not exceed a threshold $\delta$.
@@
However, as adversarial defense techniques improve, small geometric perturbations alone increasingly fail to generate adversarial examples that reliably fool classifiers — particularly when strong transferability is required [2, 38, 45]. This shortcoming has led researchers to explore unrestricted adversarial attacks, which involve larger geometric perturbations. Although “unrestricted” implies that adversarial examples need not be bounded by strict geometric distance, these examples must still remain faithful to the semantics of the original image; otherwise, the core goal of preserving the input’s meaning is lost.
@@
Zhang et al. [49] introduced a probabilistic perspective on adversarial attacks, showing that traditional geometric constraints can be interpreted as specific “distance” distributions $p_{\text{dis}}$. Generating adversarial examples then amounts to sampling from the product of $p_{\text{dis}}$ and the “victim” distribution $p_{\text{vic}}$ (representing the target classifier under attack). This approach, depicted on the left side of Figure 1, indicates that the distance distribution $p_{\text{dis}}$ need not stem solely from geometric constraints; instead, one can fit a probabilistic generative model (PGM) around the original image so that the PGM’s semantic understanding implicitly defines a semantics-based distance.
@@
Building on this probabilistic perspective, we expand the distance distribution from operating on a single image to operating on an entire concept representing an object. As shown on the right side of Figure 1, this generalization introduces a new class of adversarial attacks. Rather than perturbing a single image, we generate a fresh image that captures the same underlying concept yet deceives the classifier. We refer to this approach as a concept-based adversarial attack. Mathematically, it remains consistent with traditional adversarial attacks when viewed under the probabilistic framework. As we demonstrate, broadening the distance distribution to concept-level information reduces its gap from the victim distribution $p_{\text{vic}}$, resulting in substantially higher attack success rates.
@@
  • Concept-based adversarial attack: We introduce a new type of adversarial attack that moves beyond single-image perturbations to a concept described by a set or a model. Under a probabilistic perspective, this new approach aligns with traditional adversarial attacks.
  • Theoretical and empirical validation: We provide both theoretical proof and experimental evidence showing that expanding the attack from a single image to an entire concept reduces the distance between $p_{\text{vic}}$ and $p_{\text{dis}}$.
@@
[…]
Figure 1: Comparison of a single-image adversarial attack (left) versus our proposed concept-based adversarial attack (right). In both cases, adversarial examples $x_{\text{adv}}$ are drawn from the product of a distance distribution $p_{\text{dis}}$ and a victim distribution $p_{\text{vic}}$ [49]. On the left, $p_{\text{dis}}$ is centered on a single image $x_{\text{ori}}$, so its overlap with $p_{\text{vic}}$ is small. Consequently, adversarial examples that successfully fool the victim classifier typically lose the original image’s meaning, whereas those that preserve the original meaning fail to deceive the classifier. In contrast, on the right, $p_{\text{dis}}$ spans the original concept $\mathcal{C}_{\text{ori}}$, greatly increasing overlap with $p_{\text{vic}}$. As a result, the generated adversarial examples both maintain the concept’s meaning and easily deceive the classifier. (A green image border indicates an example that successfully fools the classifier; red indicates failure.)
@@ 2.1 Probabilistic Generative Models (PGMs) and Their Likelihoods
The goal of probabilistic generative models is to build a parameterized model $p_{\theta}$ that captures the true distribution $p$. However, in practice, we only have access to a finite dataset $\mathcal{D}=\{x_{1},\dots,x_{n}\}$, where $n$ is the size of the dataset. A common approach is to maximize the likelihood of this dataset.
For image modeling, popular models such as Variational Autoencoders (VAEs) [20] and diffusion models [14, 37] typically do not maximize the likelihood directly. Instead, they maximize a lower bound on the log-likelihood (often referred to as the evidence lower bound, or ELBO). Given an image $x$, the corresponding likelihood $p_{\theta}(x)$ can thus be estimated via this lower bound. In practice, the process of likelihood evaluation involves calculating this ELBO to estimate the likelihood [4, 27].
@@ 2.2 Adversarial Attack
where $\mathcal{D}$ measures the distance (similarity) between $x_{\text{ori}}$ and $x_{\text{adv}}$, typically via an $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, or $\mathcal{L}_{\infty}$ norm. Directly solving this constrained optimization can be challenging. To address this, Szegedy et al. [40] propose a relaxation:
@@
where $c$ is a constant, and $f$ is an objective function that guides the classifier’s predictions toward the target label. In [40], $f$ is taken to be the cross-entropy loss; Carlini and Wagner [5] present additional choices for $f$.
@@ 2.3 Probabilistic Adversarial Attack
By employing Langevin Dynamics as an optimizer for (1), Zhang et al. [49] derive a probabilistic perspective on adversarial attacks. They introduce the adversarial distribution:
@@
where $p_{\text{vic}}(x_{\text{adv}}\mid y_{\text{tar}})\propto\exp\bigl{(}-c\,f(x_{\text{adv}},y_{\text{tar}})\bigr{)}$ is the “victim” distribution emphasizing misclassification toward $y_{\text{tar}}$, and $p_{\text{dis}}(x_{\text{adv}}\mid x_{\text{ori}})\propto\exp\bigl{(}-\mathcal{D}(x_{\text{ori}},x_{\text{adv}})\bigr{)}$ is the “distance” distribution around $x_{\text{ori}}$. This formulation leverages the fact that Langevin Dynamics converges to the corresponding Gibbs distribution [23], thereby providing a probabilistic interpretation of adversarial attack generation.
@@
Zhang et al. [49] indicate that the distance distribution $p_{\text{dis}}$ can be any distribution centered around $x_{\text{ori}}$, meaning the choice of $p_{\text{dis}}$ implicitly defines the distance $\mathcal{D}$. Consequently, using a PGM centered on $x_{\text{ori}}$ as $p_{\text{dis}}$ yields a semantic-aware notion of distance. By then sampling from the corresponding adversarial distribution $p_{\text{adv}}$, one can generate semantic-aware adversarial examples.
@@ 3 Concept-based Adversarial Attack
Building on the probabilistic perspective of adversarial attacks, we can generalize the notion of an original image $x_{\text{ori}}$ to an original concept $\mathcal{C}_{\text{ori}}$. For example, $\mathcal{C}_{\text{ori}}$ could represent a specific object or identity, such as the small Corgi (dog) with slightly elongated ears in Figure 1. Rather than applying imperceptible perturbations to $x_{\text{ori}}$, concept-based adversarial attacks aim to generate an adversarial example $x_{\text{adv}}$ that preserves the conceptual identity of $\mathcal{C}_{\text{ori}}$ (i.e., the same Corgi) while allowing variation in pose, background, and viewpoint.
Technically, defining an explicit distance $\mathcal{D}(x_{\text{adv}},\mathcal{C}_{\text{ori}})$ is challenging. However, the probabilistic perspective offers a natural alternative: we define a distance distribution $p_{\text{dis}}(\cdot\mid\mathcal{C}_{\text{ori}})$ that implicitly imposes a distance measure. Formally,
@@
where $p_{\text{adv}}(\cdot\mid\mathcal{C}_{\text{ori}},y_{\text{tar}})$ is the adversarial distribution relative to the concept $\mathcal{C}_{\text{ori}}$ and the target label $y_{\text{tar}}$. The distance distribution $p_{\text{dis}}(\cdot\mid\mathcal{C}_{\text{ori}})$ is a distribution centered around the concept $\mathcal{C}_{\text{ori}}$. From a practical standpoint, one feasible approach is to construct a PGM for $\mathcal{C}_{\text{ori}}$. In practice, this is often done by fine-tuning an existing PGM on a set of images that represent the concept. Consequently, throughout the remainder of this work, we will represent the concept $\mathcal{C}_{\text{ori}}$ using a set of images $\mathcal{C}_{\text{ori}}=\{x_{\text{ori}}^{(1)},\dots,x_{\text{ori}}^{(K)}\}$, where $K$ is the number of images depicting $\mathcal{C}_{\text{ori}}$. The probabilistic adversarial attack, that manipulates only one image, can then be expressed as a special case where $\mathcal{C}_{\text{ori}}=\{x_{\text{ori}}\}$ or $|\mathcal{C}_{\text{ori}}|=1$.
@@ 3.1 Concept-based Adversarial Attacks Generate Higher Quality Adversarial Examples
From the probabilistic perspective, generating adversarial examples amounts to sampling from the overlap between $p_{\text{vic}}$ and $p_{\text{dis}}$, since $p_{\text{adv}}$ is proportional to their product [13]. For the common case of attacking a single original image $x_{\text{ori}}$, this procedure is illustrated on the left side of Figure 1. Empirical research has shown that modern robust classifiers can produce high-quality images of the target classes [33, 49, 52], causing $p_{\text{vic}}$ to concentrate on the semantics of those classes. Consequently, as $x_{\text{ori}}$ does not depict the target class, the intersection between $p_{\text{vic}}$ and $p_{\text{dis}}$ is small. Since high-quality images rarely appear in low-density regions of the distribution, the resulting adversarial examples drawn from this limited intersection tend to be of lower quality.
@@ 3.2 The Distance Between Distributions: A Theoretical Study
In the whitebox adversarial attack scenario, both the victim classifier and target label are provided, which means $p_{\text{vic}}$ remains fixed. We assume that the distance distributions $p_{\text{dis}}(\cdot\mid x_{\text{ori}})$ and $p_{\text{dis}}(\cdot\mid\mathcal{C}_{\text{ori}})$ follow Gaussian distributions with identical means $\mu$, but with $p_{\text{dis}}(\cdot\mid\mathcal{C}_{\text{ori}})$ exhibiting greater variance than $p_{\text{dis}}(\cdot\mid x_{\text{ori}})$. This assumption is reasonable since $x_{\text{ori}}$ represents an instance of concept $\mathcal{C}_{\text{ori}}$, placing their means in proximity to each other. Additionally, the distance distribution surrounding the concept $\mathcal{C}_{\text{ori}}$ should naturally demonstrate higher variance compared to that of the specific instance $x_{\text{ori}}$. The following theorem enables us to demonstrate that concept-based adversarial attacks decrease the KL divergence between $p_{\text{vic}}$ and $p_{\text{dis}}$.
@@
Let $p$ be a probability distribution and $q$ be a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. Then $KL(p||q)$ is a decreasing function of $\sigma^{2}$ when $\sigma^{2}<\mathbb{E}_{X\sim p}[(X-\mu)^{2}]$.
The proof for Theorem 1 is provided in Appendix A. By Theorem 1, it follows that $KL(p_{\text{vic}}\,\|\,p_{\text{dis}})$ decreases when $\sigma^{2}<\mathbb{E}_{X\sim p_{\text{vic}}}[(X-\mu)^{2}]$, where $\mu$ and $\sigma^{2}$ are the mean and variance of $p_{\text{dis}}$. In the probabilistic adversarial attack framework, this condition is always satisfied because it is equivalent to
$\mathbb{E}_{X\sim p_{\text{dis}}}\bigl{[}(X-\mu)^{2}\bigr{]}<\mathbb{E}_{X\sim p_{\text{vic}}}\bigl{[}(X-\mu)^{2}\bigr{]}.$
Intuitively, samples drawn from $p_{\text{vic}}$ lie farther from the mean of $p_{\text{dis}}$ than samples from $p_{\text{dis}}$ itself. If that were not the case, our core assumption that $p_{\text{dis}}$ represents a distance distribution capturing points close to $x_{\text{ori}}$ or $\mathcal{C}_{\text{ori}}$ would be violated.

### v2

# arXiv:2507.02965v2[cs.CV] 26 Feb 2026

@@ Abstract
We propose a concept-based adversarial attack framework that extends beyond single-image perturbations by adopting a probabilistic perspective. Rather than modifying a single image, our method operates on an entire concept — represented by a distribution — to generate diverse adversarial examples. Preserving the concept is essential, as it ensures that the resulting adversarial images remain identifiable as instances of the original underlying category or identity. By sampling from this concept-based adversarial distribution, we generate images that maintain the original concept but vary in pose, viewpoint, or background, thereby misleading the classifier. Mathematically, this framework remains consistent with traditional adversarial attacks in a principled manner. Our theoretical and empirical results demonstrate that concept-based adversarial attacks yield more diverse adversarial examples and effectively preserve the underlying concept, while achieving higher attack efficiency. Code and examples can be found at https://github.com/andiac/ConceptAdv.
@@ 1 Introduction
Adversarial attacks aim to deceive a classifier while preserving the original meaning of the input object (Dalvi et al., 2004; Lowd and Meek, 2005a; b; Biggio and Roli, 2018). We refer to the manipulated instance as an adversarial example. Early work by Szegedy et al. (2014) and Goodfellow et al. (2015) introduced adversarial attacks against deep learning models for images.
@@
In the image-based adversarial setting, it is widely accepted that controlling the geometric distance between the adversarial example and its original image is crucial for maintaining the original image’s meaning. Consequently, many adversarial attack algorithms constrain perturbations using norms such as $L_{1}$, $L_{2}$, or $L_{\infty}$. Moreover, given the rapid progress in machine learning research, fair comparisons across different adversarial methods have become essential. Numerous benchmarks and competitions (Madry et al., 2018; Croce et al., 2020; Dong et al., 2021) therefore focus on attack success rates under the constraint that geometric distance does not exceed a threshold $\delta$.
@@
However, as adversarial defense techniques improve, small geometric perturbations alone increasingly fail to generate adversarial examples that reliably fool classifiers, particularly when strong transferability is required (Song et al., 2018; Xiao et al., 2018b; Bhattad et al., 2020). This shortcoming has led researchers to explore unrestricted adversarial attacks, which involve larger geometric perturbations. Although “unrestricted” implies that adversarial examples need not be bounded by strict geometric distance, these examples must still remain faithful to the semantics of the original image; otherwise, the core goal of preserving the input’s meaning is lost.
@@
Zhang et al. (2024b) introduced a probabilistic perspective on adversarial attacks, demonstrating that traditional geometric constraints can be interpreted as specific “distance” distributions $p_{\text{dis}}$. Under this view, generating adversarial examples amounts to sampling from the product of $p_{\text{dis}}$ and the “victim” distribution $p_{\text{vic}}$, which represents the target classifier under attack. Importantly, $p_{\text{dis}}$ need not be induced solely by geometric distance. Instead, one can fit a probabilistic generative model (PGM) around the original image, allowing the PGM’s semantic representation to implicitly define a semantics-based notion of distance. As illustrated on the left side of Figure 1, Zhang et al. (2024b) indicate that $p_{\text{dis}}$ should be centered on the original image.
@@
Building on Zhang et al. (2024b)’s probabilistic perspective, we expand the distance distribution $p_{\text{dis}}$ from operating on a single image to operating on an entire concept, which can be represented by a probability distribution over images that correspond to the same underlying object, identity or category. As shown on the right side of Figure 1, this generalization introduces a new class of adversarial attacks. Rather than perturbing a single image, we generate a fresh image that captures the same underlying concept yet deceives the classifier. We refer to this approach as a concept-based adversarial attack. Mathematically, it remains consistent with traditional adversarial attacks when viewed under the probabilistic framework. As we demonstrate, broadening the distance distribution to concept-level information reduces its gap from the victim distribution $p_{\text{vic}}$, resulting in substantially higher attack success rates.
[…]
Figure 1: Comparison of a single‐image adversarial attack (left) versus our proposed concept‐based adversarial attack (right). In both cases, adversarial examples $x_{\text{adv}}$ are drawn from the product of a distance distribution $p_{\text{dis}}$ and a victim distribution $p_{\text{vic}}$. On the left, $p_{\text{dis}}$ is centered on a single image $x_{\text{ori}}$, so its overlap with $p_{\text{vic}}$ is small. Consequently, adversarial examples that successfully fool the victim classifier typically lose the original image’s meaning, whereas those that preserve the original meaning fail to deceive the classifier. In contrast, on the right, $p_{\text{dis}}$ spans the original concept $\mathcal{C}_{\text{ori}}$, greatly increasing overlap with $p_{\text{vic}}$. As a result, the generated adversarial examples both maintain the concept’s meaning and easily deceive the classifier. (A green image border indicates an example that successfully fools the classifier; red indicates failure.)
@@
  • Concept-based adversarial attack: We introduce a new type of adversarial attack that moves beyond single-image perturbations to a concept represented by a distribution, this new approach aligns with traditional adversarial attacks in a principled manner.
@@
  • Theoretical and empirical validation: We provide both theoretical proof and experimental evidence showing that expanding the attack from a single image to an entire concept reduces the distance between $p_{\text{vic}}$ and $p_{\text{dis}}$, boosting the attack efficiency.
@@
  • Practical Guidelines and Scenarios: We provide practical guidelines and example application scenarios, detailed in Appendix K.
@@ 2.1 Probabilistic Generative Models (PGMs) and Their Likelihoods
The goal of probabilistic generative models is to learn a parameterized distribution $p_{\theta}$ that approximates the true distribution $p$. In practice, we only observe a finite dataset $\mathcal{D}=\{x_{1},\dots,x_{n}\}$, and training is typically done by maximizing its likelihood. For image modeling, popular approaches such as VAEs (Kingma et al., 2013) and diffusion models (Song and Ermon, 2019; Ho et al., 2020) optimize a lower bound on the log-likelihood (the ELBO) rather than the likelihood itself. Thus, likelihood estimation in practice amounts to computing this ELBO (Burda et al., 2015; Nalisnick et al., 2019).
@@ 2.2 Adversarial Attack
where $\mathcal{D}$ measures the distance (similarity) between $x_{\text{ori}}$ and $x_{\text{adv}}$, typically via an $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, or $\mathcal{L}_{\infty}$ norm. Directly solving this constrained optimization can be challenging. To address this, Szegedy et al. (2014) propose a relaxation:
@@
where $c$ is a constant, and $f$ is an objective function that guides the classifier’s predictions toward the target label. In Szegedy et al. (2014)’s work, $f$ is taken to be the cross-entropy loss; Carlini and Wagner (2017) present additional choices for $f$.
@@ 2.3 Probabilistic Adversarial Attack
By employing Langevin Dynamics as an optimizer for equation 1, Zhang et al. (2024b) derive a probabilistic perspective on adversarial attacks. They introduce the adversarial distribution:
@@
where $p_{\text{vic}}(x_{\text{adv}}\mid y_{\text{tar}})\propto\exp\bigl(-c\,f(x_{\text{adv}},y_{\text{tar}})\bigr)$ is the “victim” distribution emphasizing misclassification toward $y_{\text{tar}}$, and $p_{\text{dis}}(x_{\text{adv}}\mid x_{\text{ori}})\propto\exp\bigl(-\mathcal{D}(x_{\text{ori}},x_{\text{adv}})\bigr)$ is the “distance” distribution around $x_{\text{ori}}$. This formulation leverages the fact that Langevin Dynamics converges to the corresponding Gibbs distribution (Lamperski, 2021), thereby providing a probabilistic interpretation of adversarial attack generation.
@@
Zhang et al. (2024b) indicate that the distance distribution $p_{\text{dis}}$ can be any distribution centered around $x_{\text{ori}}$, meaning the choice of $p_{\text{dis}}$ implicitly defines the distance $\mathcal{D}$. Consequently, using a PGM centered on $x_{\text{ori}}$ as $p_{\text{dis}}$ yields a semantic-aware notion of distance. By then sampling from the corresponding adversarial distribution $p_{\text{adv}}$, one can generate semantic-aware adversarial examples.
@@ 3 Concept-based Adversarial Attack
3.1 Concept Distribution
We aim to extend adversarial attacks from operating on a single original image to operating on an original concept $\mathcal{C}_{\text{ori}}$. A concept is inherently abstract and subjective: it may refer to a specific physical object (e.g., a rubber duck), a particular identity such as the long-eared corgi puppy with a lighter left cheek shown in Figure 1, or a broader class such as “corgi,” regardless of age, size, or specific attributes. Although defining a concept in an absolute sense is difficult (Poeta et al., 2023), we can represent it through a concept distribution, denoted by $p(\cdot\mid\mathcal{C}_{\text{ori}})$.
This distribution serves as an interface through which users can specify what the concept is. In practice, we recommend two ways for users to instantiate their notion of a concept:
  • Direct specification of a concept distribution: The user may already possess a generative model or any other mechanism that directly provides a distribution $p(\cdot\mid\mathcal{C}_{\text{ori}})$ representing the concept.
  • Constructing the distribution from an image set: The user may collect a set of images depicting the desired concept (e.g., different poses of the same corgi in Figure 1), and then train or fine-tune a probabilistic generative model (PGM) on this set to obtain the corresponding concept distribution $p(\cdot\mid\mathcal{C}_{\text{ori}})$. Here, $\mathcal{C}_{\text{ori}}$ is a set of images $\mathcal{C}_{\text{ori}}=\{x_{\text{ori}}^{(1)},\dots,x_{\text{ori}}^{(K)}\}$, where $K$ is the number of images depicting $\mathcal{C}_{\text{ori}}$. .
In the remainder of this paper, we demonstrate the second approach, as it allows us to clearly showcase how concept-level information can be incorporated into adversarial attacks using accessible image data and standard generative modeling pipelines.
3.2 Concept Distribution as a Distance distribution
Building on the probabilistic perspective of adversarial attacks (Zhang et al., 2024b), a distribution used as a distance distribution can implicitly define a notion of distance. Therefore, by using the concept distribution defined in the previous subsection as the distance distribution, we implicitly define the distance between an adversarial example and the underlying concept. Formally,
@@ 3.2 Concept Distribution as a Distance distribution
where $p_{\text{adv}}(\cdot\mid\mathcal{C}_{\text{ori}},y_{\text{tar}})$ is the adversarial distribution relative to the concept $\mathcal{C}_{\text{ori}}$ and the target label $y_{\text{tar}}$. The distance distribution $p_{\text{dis}}(\cdot\mid\mathcal{C}_{\text{ori}})$ is the concept distribution.
Comparing (2) and (3) shows that the only modification is replacing $x_{\text{ori}}$ with $\mathcal{C}_{\text{ori}}$. Hence, the probabilistic adversarial attack (Zhang et al., 2024b) is the special case $\mathcal{C}_{\text{ori}}=\{x_{\text{ori}}\}$ (i.e., $|\mathcal{C}_{\text{ori}}|=1$). This straightforward and compact expansion allows us to heavily reuse the implementation of the probabilistic adversarial attack, making probabilistic adversarial attack (ProbAttack [^1] Both Zhang et al. (2024b) and our work present a methodology applicable to any PGM (e.g., VAE, energy-based, or diffusion). Since diffusion models are the most powerful PGMs, we adopt them throughout this paper. We use ProbAttack to denote the diffusion-based implementation of Zhang et al. (2024b).) a natural ablation baseline for our method.
While intuition suggests that expanding the perturbation space should produce stronger adversarial examples, rigorous justification is needed. In the following sections, we adopt the probabilistic perspective, presenting both theoretical analysis and empirical evidence to demonstrate how this expansion enhances attack effectiveness without compromising perceptual quality.
@@ 3.3 Concept-based Adversarial Attacks Generate Higher Quality Adversarial Examples
From the probabilistic perspective, generating adversarial examples amounts to sampling from the overlap between $p_{\text{vic}}$ and $p_{\text{dis}}$, since $p_{\text{adv}}$ is proportional to their product (Hinton, 2002). For the common case of attacking a single original image $x_{\text{ori}}$, this procedure is illustrated on the left side of Figure 1. Empirical research has shown that modern robust classifiers can produce high-quality images of the target classes (Santurkar et al., 2019; Zhang et al., 2024b; Zhu et al., 2021), causing $p_{\text{vic}}$ to concentrate on the semantics of those classes. Consequently, as $x_{\text{ori}}$ does not depict the target class, the intersection between $p_{\text{vic}}$ and $p_{\text{dis}}$ is small. Since high-quality images rarely appear in low-density regions of the distribution, the resulting adversarial examples drawn from this limited intersection tend to be of lower quality.
@@ 3.4 The Distance Between Distributions: A Theoretical Study
In the whitebox adversarial attack scenario, both the victim classifier and target label are provided, which means $p_{\text{vic}}$ remains fixed. Let $p_{\text{dis}}(\cdot\mid\mathcal{C}_{\text{ori}})$ be a Gibbs distribution of the form $p_{\text{dis}}(x\mid\mathcal{C}_{\text{ori}})\propto\exp(-\beta D(x,\mu))$, where $D$ is a distance function measuring the discrepancy between a point $x$ and the concept center $\mu$. The following theorem shows that, under suitable conditions, increasing the dispersion of $p_{\text{dis}}$ (i.e., decreasing $\beta$ [^2] A smaller $\beta$ corresponds to a higher “temperature” in the Gibbs distribution, which makes $p_{\text{dis}}$ more dispersed.) reduces the KL divergence between $p_{\text{vic}}$ and $p_{\text{dis}}$:
@@
Let $p$ be a probability distribution and $q$ be a Gibbs distribution of the form
$q(x)=\frac{\exp(-\beta D(x,\mu))}{Z(\beta)},$
where $Z(\beta)$ is the normalizing constant, $\mu$ is a constant and $D$ is a distance function. Then $KL(p\,\|\,q)$ is a increasing function of $\beta$ whenever $\mathbb{E}_{X\sim p}[D(X,\mu)]>\mathbb{E}_{X\sim q}[D(X,\mu)]$.
The proof is provided in Appendix A. By Theorem 1, we see that $KL(p_{\text{vic}}\,\|\,p_{\text{dis}})$ decreases as $\beta$ decreases, provided that $\mathbb{E}_{X\sim p}[D(X,\mu)]>\mathbb{E}_{X\sim q}[D(X,\mu)]$. In the probabilistic adversarial attack framework, this condition is always satisfied because samples drawn from $p_{\text{vic}}$ lie farther from the mean of $p_{\text{dis}}$ than samples drawn from $p_{\text{dis}}$ itself. If this were not the case, the fundamental assumption that $p_{\text{dis}}$ represents a distance distribution concentrated around $x_{\text{ori}}$ or $\mathcal{C}_{\text{ori}}$ would be violated.
In practice, different PGMs may be used to model $p_{\text{dis}}$. When an energy-based model (EBM) is adopted, it explicitly learns a Gibbs distribution (LeCun et al., 2006). When a diffusion model (via score matching) is used, it instead learns an implicit representation of the corresponding energy function (Song and Ermon, 2019). Consequently, treating $p_{\text{dis}}$ as a Gibbs distribution in this section is fully consistent with practical implementations, and is also aligned with the probabilistic adversarial attack formulation in Section 2.3.